# An Informative Planning Framework for Target Tracking and Active Mapping in Dynamic Environments with ASVs

### v2

@@ VII Conclusions and Future Work
In this study, we introduce an IPP framework to actively map moving targets in dynamic environments. We introduce a spatiotemporal prediction network to estimate uncertain target position distributions over time due to drift, and a new IPP utility for target tracking. We use a dynamic occupancy mapping approach to maintain an accurate map. Simulation results show that our IPP approach with the proposed utility improves target tracking by redetecting targets more frequently, as compared to existing methods that use only entropy as the utility. The ablation studies further highlight the importance of our spatiotemporal network and our dynamic mapping method. Finally, field experiments with an ASV demonstrate our method’s effectiveness in real-world monitoring scenarios. Although our proposed IPP approach performs well in both simulations and real experiments, the replanning step is computationally expensive, limiting the decision frequency. Future work will investigate learning-based techniques to develop efficient and faster IPP methods for monitoring in dynamic environments.

### v3

@@ VII Conclusions and Future Work
In this study, we introduce an IPP framework to actively map moving targets in dynamic environments. We introduce a spatiotemporal prediction network to estimate uncertain target position distributions over time due to drift, and a new IPP utility for target tracking. We use a dynamic occupancy mapping approach to maintain an accurate map. Simulation results show that our proposed utility improves target tracking by redetecting targets more frequently, as compared to existing methods that use only entropy as the utility. Ablation studies further highlight the importance of our spatiotemporal network and our dynamic mapping method. Finally, field experiments with an ASV demonstrate the validity of our framework in real-world monitoring scenarios. While our approach can track targets efficiently in real-time, it relies on wind measurements. Therefore, estimating the drift of floating objects purely based on observations is an avenue for further work. Another direction for future work is multi-robot spatiotemporal monitoring.
Acknowledgments
The authors thank Joel Jose, Jon Urcelay, and Elisha Adiburo for assistance with the field experiments.